# leftmargin = 20pt, rightmargin = 20pt, backgroundcolor = bggray, middlelinecolor = black, roundcorner = 15

### v1

@@ 7.2 Service provider
Algorithm 4 is an implementation of the routine for the service providers. For announcement and advertisement of their service, the providers publish a discoverability event of kind 31990 (Event Type 5). Initially, the service providers wait until they receive a job request. Then they signal initiation through a job feedback event. Upon receiving any required payments they proceed with the optimization routine, as the customer has requested. The optimization procedure may be one of the run options (FEDAVG or DiLoCo) and the task for either run option may be Inner or Outer optimization. Depending on the run option the service providers calls the corresponding algorithmic routine (Algorithm 5 or Algorithm 6). Periodically the service provider announces the status of the job through a Job Feedback event and provides a progress of the computation, for instance see Algorithm 5. After the completion of the optimization routine the service providers respond with a job result event, and they expect any pending payment to be settled. If the customer does not complete payment(s) after a reasonable time period despite the delivery of accurate outputs, then the service providers may stop receiving job request from the specific customer immediately.
Federated Averaging vs Distributed Low-Communication Training of LLMs (DiLoCo)
@@
The federated averaging algorithm [34] (FEDAVG) is a distributed and communication efficient learning method of AI models from decentralized data. This decentralized computation is known as federated learning; a set of different devices learn a shared model by aggregating locally-computed updates. In our design we consider as one optimization approach the FEDAVG. The customer provides a common model specification and the data to the servicce providers. The service providers have the role of different devices, they obtain a subset of the data-set and they train the shared model. Upon completion of the computation they return the optimized model parameters, and they receive a payment.
@@
Although the FEDAVG is a successful approach for training neural netowrks and AI in a decentralized fashion, we also consider a variant of FEDAVG that has numerous advantages over the classical approach. Specifically, the DiLoCo algorithm [8] enables training of LLMs under constrained communication. In fact, the inner part of optimization (jobs by Service providers) requires the greatest amount of computation, while the request for a new job is being decreased significantly. This provides additional flexibility to our design since it minimizes the number of job requests by the customer. Nevertheless, the service providers have the option to frequently provide information (e.g. job status, partial output) to the customer through the Job Feedback events. We proceed by presenting the implementation of the Inner Optimization and Outer Optimization routines for both FEDAVG and DiLoCo algorithms.
@@
For the Outer Optimization routine, the customer has the option to run the outer optimization, or assign it to a service provider. Then the customer (or a service provider) runs the outer part of FEDAVG (example aggregation step) or DiLoCo [8, 40, Nesterov Momentum]. The run option is determined by the customer. The algorithm returns the updated model parameter.
@@ 8 Conclusion
The NOSTR protocol is a relatively new paradigm shift for reliable and decentralized social web applications. In this work we introduced a design and protocol flow for LLM and AI training on decentralized marketplace system, that builds upon the NOSTR. The fast communication though websockets and the flexibility of development of the protocol enables the potential for a competitive, open and censorship resistant computation network, at which users (customers) provide data and a payment, and they receive a trained AI model. The integrated payment protocol flow rules and the lightning network offer fast, reliable and with zero transaction fee payments without a trusted party. As a result, service providers can maximize their profit and compete within an open market. Further, this payment system supports microtranscactions with instant settlement, which is essentially useful as anti-spam mechanism, reliable form of handshaking between customers and service providers, on the fly payments upon receiving partial results of the computation, or upon validation of the final output.
We considered a simple but effective mechanism for validation of the output, by periodically checking the progress of the trained model through a validation dataset. However, we believe that other alternative solutions for verifying accuracy of the computation may be considered as well. Such robust approaches can be designed by considering smart contracts and the BitVM mechanism [41], which enables the possibility of any computable function to be verified. Inclusion of such mechanisms for verification, together with a reputation system, or application specific protocol flow variations for alternative AI algorithms such as DeepSpeed [42] & Zero Redundancy Optimizer [43] or distributed LLM deployment and inference [44, 45, 46, 47, 48, 49, 50, 51] could be potentially considered tasks for future work.

### v2

@@ 7.2 Service Provider
Algorithm 4 implements the routine for service providers. To advertise their services, service providers publish a discoverability event of kind 31990 (Event Type 5). Initially, the service providers wait until they receive a job request. Then they signal readiness through a job feedback event. Upon receiving any required payments, they execute the optimization routine, as requested by the customer. The optimization procedure uses one of two run options (FEDAVG or DiLoCo) and the task may be either Inner or Outer optimization. Depending on the run option the service providers calls the corresponding algorithmic routine (Algorithm 5 or Algorithm 6). Periodically, the service provider announces the status of the job through a Job Feedback event and provides progress updates on the computation, for instance see Algorithm 5. After completing the optimization routine, service providers respond with a job result event, and expect any pending payment to be settled. If the customer does not complete payment(s) after a reasonable time period despite the delivery of accurate outputs, service providers may immediately stop accepting job requests from that customer.
Algorithm 4 Service Provider
1:SP_NPUB: service provider public key, <computation information & time>, <lightning address>
2:Connection through websockets to all relays in RELAYS Where events are published at
3:Publish an event (NIP-89) of kind for discoverability Event Type 5
4:while TRUE do
5: Routine: Search for JOB REQUEST for provider SP_NPUB
6: if JOB REQUEST with appropriate KIND was found then Event Type 2
7: Publish Job Feedback (KIND: 7000) with STATUS = payment_request Event Type 4
8: Wait until the initial payment is submitted, fetch and validate Payment Receipt (KIND: 9735)
9: if payment = successful then
10: Publish Job Feedback (KIND: 7000) with STATUS = processing Event Type 4
11: else
12: Publish Job Feedback (KIND: 7000) with STATUS = error
13: Fetch model_parameters, DATA_SOURCES(SP_NPUB), RUN_OPTION, TASK from JOB REQUEST
14: if RUN_OPTION = FEDAVG and TASK = INNER then
15: Call Inner_Optimization( model_parameters, DATA_SOURCES , FEDAVG )
16: else if RUN_OPTION = FEDAVG and TASK = OUTER then
17: Call Outer_Optimization( model_parameters, DATA_SOURCES , FEDAVG )
18: else if RUN_OPTION = DiLoCo and TASK = INNER then
19: Call Inner_Optimization( model_parameters, DATA_SOURCES , DiLoCo )
20: else if RUN_OPTION = DiLoCo and TASK = OUTER then
21: Call Outer_Optimization( model_parameters, DATA_SOURCES , DiLoCo )
22: Publish Job Feedback (KIND: 7000) with STATUS = success Event Type 4
23: Publish a JOB RESULT event with "output" the result of the optimization Event Type 3
Federated Averaging vs Distributed Low-Communication
Training of LLMs (DiLoCo)
@@
The federated averaging algorithm [34] (FEDAVG) is a distributed and communication-efficient learning method for AI models trained on decentralized data. This decentralized computation is known as federated learning; a set of different devices learn a shared model by aggregating locally-computed updates. In our design, we consider the FEDAVG algorithm as one optimization approach. The customer provides a common model specification and the data to the service providers. The service providers have the role of different devices; they obtain a subset of the dataset and they train the shared model. Upon completion of the computation, they return the optimized model parameters and receive payment.
@@
Although the FEDAVG is a successful approach for training neural networks and AI in a decentralized fashion, we also consider a variant that has numerous advantages over the classical approach. Specifically, the DiLoCo algorithm [8] enables the training of LLMs under constrained communication. In fact, the inner part of optimization (jobs by service providers) requires the greatest amount of computation, while the frequency of new job requests is significantly reduced. This provides additional flexibility to our design by minimizing the number of job requests from the customer. Nevertheless, service providers can frequently provide information (e.g. job status, partial output) to the customer through the Job Feedback events. We proceed by presenting the implementation of the Inner Optimization and Outer Optimization routines for both FEDAVG and DiLoCo algorithms.
@@
Algorithm 5 Inner_Optimization
1:RUN_OPTION: <FEDAVG or DiLoCo>, DATA: DATA_SOURCES, model_parameters:
2:Connection through websockets to all relays in RELAYS Where events are published at
3:
4:if RUN_OPTION = FEDAVG then Run the inner part of FEDAVG
5: data DATA_SOURCES(SP_NPUB)
6: Split data into batches
7: for epoch do
8: Publish Job Feedback (KIND: 7000) with STATUS = processing Event Type 4
9: for batch do
10:
11:else if RUN_OPTION = DiLoCo then
12: for epoch do
13: Sample a batch of data from DATA_SOURCES(SP_NPUB)
14: Run AdamW() AdamW [39, Algorithm 2]
15: Periodically publish Job Feedback (KIND: 7000) with STATUS = processing Event Type 4
16: AdamW()
17:return Returns
@@
For the Outer Optimization routine, the customer can either run the outer optimization locally or assign it to a service provider. Then the customer (or a service provider) runs the outer part of FEDAVG (e.g., aggregation step) or DiLoCo [8, 40, Nesterov Momentum]. The run option is determined by the customer. The algorithm returns the updated model parameters.
Algorithm 6 Outer_Optimization
1:RUN_OPTION: <FEDAVG or DiLoCo>, DATA: DATA_SOURCES, model_parameters: , the last "outputs" of the Inner Optimization jobs:
2:Connection through websockets to all relays in RELAYS Where events are published at
3:if RUN_OPTION = FEDAVG then Run the aggregation (outer) part of FEDAVG
4: For some weights
5:else if RUN_OPTION = DiLoCo then Run the Outer part of DiLoCo [8]
6:
7: Run Nesterov Momentum: See FedMom [40, Algorithm 3]
8: (Periodically) publish Job Feedback (KIND: 7000) with STATUS = processing Event Type 4
9: Nesterov_Momentum ( , )
10:return
8 Proof-of-Concept Implementation
To validate the FEDSTR protocol design, we developed a working proof-of-concept implementation demonstrating federated learning over the NOSTR network with cryptographic integrity guarantees. The primary contribution of this implementation is the demonstration that (1) the NOSTR protocol can successfully orchestrate distributed federated learning, (2) model validation ensures model integrity across untrusted relays, and (3) the system operates on public, decentralized infrastructure without trusted intermediaries.
8.1 Implementation Overview
The proof-of-concept implements the complete FEDSTR protocol flow as described in Section 7, including:
Customer coordinator managing federated learning rounds
Data Vending Machines (DVMs) as federated learning service providers
Communication over public NOSTR relays: All events are cryptographically signed using Schnorr signatures and propagated through multiple public NOSTR relays, demonstrating the protocol’s decentralized nature.
SHA-256 cryptographic hash verification for model integrity at every exchange
Validation of DVM outputs
Model storage with hash-based content addressing
Note: Similarly to exchanging notes (messages on NOSTR), exchanging model parameters also satisfies the following:
Relays cannot modify model parameters without detection
Man-in-the-middle attacks are cryptographically prevented
The protocol achieves model integrity without trusted intermediaries
Hash verification adds negligible computational overhead
Code Availability: The proof-of-concept is open source and available on GitHub [41].
8.2 Storage Architecture
The implementation supports multiple storage backends for model parameters.
Current Implementation: Local filesystem storage (file:// URLs) with SHA-256 hash-based filenames. Models are serialized to raw bytes (400KB) and stored at /tmp/fedstr_models/model_<hash>.bin.
Example of Storage Operations:
Upload: DVM serializes model computes hash saves to storage publishes URL + hash
Download: Customer retrieves URL from JobResult downloads bytes verifies hash deserializes model
Extensibility: The storage layer is abstracted to support future backends:
HTTP servers for remote storage (Tested already with private storage service)
IPFS for content-addressed distributed storage
Blossom servers (NOSTR-native file storage)
NIP-94 file metadata events (see Section 8.4)
All the coordination and ML operations are storage-agnostic – integrity is guaranteed regardless of storage backend, as long as the model parameters’ hash matches.
8.3 Current Limitations
The proof-of-concept demonstrates core protocol functionality but intentionally simplifies certain components for clarity:
Payments: Dummy Lightning Network payment implementation (no actual bitcoin transfers). Events include bolt11 invoice fields, but payment verification is stubbed. Full NIP-57 integration is planned.
DVM Identity: DVMs generate fresh keypairs on startup. Production systems should use persistent keypairs stored securely.
Deployment: Single-machine simulation (DVMs run as separate processes). Protocol is designed for multi-machine deployment, which remains to be tested at scale.
Reputation: No reputation system implemented. Trust relies solely on cryptographic verification and validation.
Storage: Local file storage and HTTP servers for remote storage. Decentralized cloud storage integration (Blossom, IPFS) is implemented but not yet tested in production.
These limitations do not affect the core protocol validation – they represent engineering work required for production deployment.
8.4 Future Work – NIP-94 Integration
While the current proof-of-concept embeds model URLs directly in JobResult events (Kind 6000), future versions could leverage NIP-94 (File Metadata) [42] for improved model discoverability and relay indexing.
8.4.1 NIP-94 Protocol Overview
NIP-94 defines a standardized event type (Kind 1063) for publishing file metadata to NOSTR relays. Files are described with tags including URL, MIME type, SHA-256 hash, size, and optional metadata (thumbnails, descriptions, etc.). Relays can index these events, enabling file discovery and categorization.
8.4.2 Potential Benefits for FEDSTR
1.
Standardization: Trained models published as standard NOSTR file events, interoperable with other NOSTR file applications.
2.
Discoverability: Relays can index trained models by dataset, accuracy, model architecture, etc., enabling "model marketplace" discovery.
3.
Rich Metadata: Attach training metrics, model architecture details, dataset information, and visualization thumbnails to model files.
4.
Reusability: Same trained model can be referenced from multiple JobResults without event duplication.
5.
Relay Indexing: Specialized "model relays" could index only Kind 1063 events, creating curated model repositories.
8.4.3 Proposed Implementation
Under a NIP-94 integration, the protocol flow would be modified as follows:
1.
DVM completes training and uploads model to storage (unchanged)
2.
DVM publishes NIP-94 file metadata event (Kind 1063):
"kind": 1063,
"tags": [
["url", "https://blossom.server/model_abc123.bin"],
["x", "abc123..."], // SHA-256 hash
["m", "application/octet-stream"],
["size", "409245"],
["alt", "FEDSTR MNIST Round 3 - 97.28% accuracy"]
"content": "Federated learning model trained on MNIST"
3.
DVM publishes JobResult (Kind 6000) referencing the NIP-94 event:
"kind": 6000,
"tags": [
["job_id", "abc123"],
["e", "<nip94_event_id>", "wss://relay.damus.io"],
["loss", "0.0875"],
["accuracy", "0.9728"]
4.
Customer retrieves JobResult, follows the event reference to NIP-94 event, downloads model from URL, verifies hash (unchanged verification protocol)
8.4.4 Recommendation
We recommend the following strategy:
Private FL Applications: Use direct URL embedding in JobResult (current approach). Simpler, faster, sufficient for closed federations.
Public Model Sharing: Use NIP-94 dual-event approach. Enables discoverability and marketplace features.
Hybrid: Publish both - embed URL in JobResult for efficiency, also publish NIP-94 for discoverability.
The current protocol design is compatible with all three approaches – the choice can be made at deployment time based on application requirements. Future versions of the PoC will include NIP-94 support as a configurable option.
8.5 Deployment Considerations
The proof-of-concept demonstrates that FEDSTR can operate on existing NOSTR infrastructure without modification:
Public Relays: Successfully used relay.damus.io and nos.lol (free, public, no registration required)
No Special Infrastructure: No dedicated "FEDSTR relays" needed – standard NOSTR relays handle all event types
Client Compatibility: DVMs and customers are standard NOSTR clients – can use existing libraries (nostr-sdk, NDK, etc.)
Incremental Deployment: Can scale to multiple DVMs without protocol changes
Self-Hosted Infrastructure (Optional): Users have the option to run their own relays and storage servers for maximum privacy and control. By self-hosting, organizations can:
Monitor all coordination and events in real-time
Guarantee data locality and regulatory compliance
Optimize network performance (reduce latency, increase bandwidth)
Maintain complete privacy (events never leave private infrastructure)
Control relay policies (access control, rate limiting, etc.)
Critically, self-hosted deployments use the exact same protocol as public deployments – no modifications required. A private FEDSTR federation can seamlessly transition to public relays or vice versa by simply changing relay URLs, demonstrating the protocol’s flexibility.
This demonstrates a key advantage of building on NOSTR: the protocol inherits the existing decentralized infrastructure without requiring specialized setup. Simultaneously, it supports fully private deployments for organizations with strict privacy or performance requirements.
8.6 Summary of Proof-of-Concept Contributions
The PoC validates the following key claims:
1.
Protocol Correctness: The FEDSTR event flow (JobRequest JobFeedback JobResult) successfully orchestrates federated learning across multiple DVMs
2.
Cryptographic Integrity: Cryptographic verification ensures integrity across untrusted relay infrastructure
3.
Decentralization: The system operates on public NOSTR relays with no trusted intermediaries
4.
Infrastructure Reuse: FEDSTR works on existing NOSTR infrastructure without modifications
These results demonstrate that decentralized federated learning is viable on the NOSTR protocol, validating the core thesis of this work. Complete implementation details and source code are available at
https://github.com/ConstantinosNikolakakis/Fedstr.
@@ 9 Conclusion
The NOSTR protocol represents a paradigm shift for reliable and decentralized social web applications. In this work, we introduced a design and protocol flow for LLM and AI training on decentralized marketplace system, which builds upon NOSTR. The fast communication through websockets and the flexibility of protocol development enables the possibility of a competitive, open, and censorship-resistant computation network. In this network, users (customers) provide data and payment, and receive a trained AI model. The integrated payment protocol and the Lightning network offer fast, reliable payments with virtually zero transaction fees and without a trusted party. As a result, service providers can maximize their profit and compete within an open market. Further, this payment system supports microtransactions with instant settlement, which serves as an effective anti-spam mechanism, provides a reliable form of handshaking between customers and service providers, enables on-the-fly payments upon receiving partial computational results, and supports payments upon validation of the final output.
We validated the FEDSTR protocol design through a working proof-of-concept
implementation demonstrating successful federated learning over public and private NOSTR
relays with cryptographic integrity guarantees. The implementation proves that:
(1) The NOSTR protocol can successfully orchestrate distributed federated
learning without trusted intermediaries,
(2) Cryptographic hash verification ensures model integrity across untrusted
relay infrastructure,
(3) The protocol operates on existing public NOSTR infrastructure without
requiring specialized setup.
These results demonstrate the viability of FEDSTR for real-world decentralized
federated learning applications, validating the core thesis that the NOSTR
protocol provides a suitable foundation for AI training, inference marketplaces, and agentic AI systems.
We considered a simple but effective mechanism for validation of the output, by periodically checking the progress of the trained model through a validation dataset. However, alternative solutions for verifying computational accuracy could also be considered. Such robust approaches can be designed using smart contracts and the BitVM mechanism [43], which enables verification of any computable function. Future work may also explore the inclusion of verification mechanisms and reputation systems, as well as application-specific protocol flow variations supporting alternative AI algorithms—such as DeepSpeed [44], Zero Redundancy Optimizer [45], and distributed LLM deployment and inference [46, 47, 48, 49, 50, 51, 52, 53].